# Estimating velocities of infectious disease spread through spatio-temporal log-Gaussian Cox point processes

### v1

@@ 5 Application
In this section, we show a real application of our method to model the spread and velocities of an infectious disease. In particular, we use the locations of COVID-19 infected individuals in Cali, one of the largest and most populated cities in Colombia. Data were provided by the Cali Municipal Public Health Secretary [Municipal Public Health Secretary of Cali,] and have confirmed locations of 208,551 infected individuals from 2020-03-01 to 2022-04-25. To begin, the geographical coordinates of the data were converted to UTM coordinates, and the values in the map are presented in meters. To model the LGCP, we construct a triangulated mesh covering the study region, in this case, the city boundary of Cali, Colombia. Once we have fitted the model, we make predictions in a grid with cells equally spaced 300 meters apart, which results in 1085 cells inside the city limits. For the temporal component, for simplicity, we use the 80 first days. In the image 7, we present the location of infected people on 4 different days in Cali, Colombia.
Figure 7: Locations of infected people in Cali, Colombia, for the days 2020-03-11, 2020-03-31, 2020-04-20 and 2020-05-10.
@@ 5.1 Intensity function
Following the ideas of Section 3.1, we apply a LGCP model with intensity function decomposed as the product of the spatial $\eta(\textbf{u})$, temporal $\mu(t)$, and a residual spatio-temporal variation $\exp\left(\zeta(\textbf{u},t)\right)$ as follows
$\Lambda(\textbf{u},t)=\eta(\textbf{u})\cdot\mu(t)\cdot\exp\left(\zeta(\textbf{u},t)\right).$ (16)
In practice, this decomposition requires first computing the spatial and temporal components separately and then using this information as an offset to compute the spatio-temporal component. The computation of each component is detailed below.
The spatial component represents the expected number of infected individuals in the city of Cali. To estimate this, we consider as a study region the city boundary of Cali. We use the R-INLA package to construct a triangulated mesh in this region, and compute the expected number of events using the dual mesh, which comprises a series of polygons surrounding each vertex of the original mesh [Moraga, 2023]. On the other hand, we obtain the population density in Cali from WorldPop, an open-access tool for population data. In this case, we used a resolution of $100\times 100$ $m$ cells defined in a square region that includes Cali (Figure 8). Then, the expected number of events in each polygon of the dual mesh is computed as the area of the polygon multiplied by the average population of the cells in that polygon [Krainski et al., 2018].
Figure 8: Population density in Cali, Colombia.
To estimate the temporal component representing the daily expected number of infected people, we use a spline model as explained in Section 3.1.1. The selected basis is related to the structure of the data. It takes into account first the day-of-week effects, periodical effects described by sine and cosine functions called Fourier basis, and finally an overall rising trend over the time captured by polynomial structure. The selected number of basis in each case was related to the simplicity and explainability of the model. In summary, the model has the following form
@@
Figure 9: Total number of daily COVID-19 cases (blue dots) are presented with the fitted model (solid red line) and the confidence interval of the prediction over time (pink hatched area).
Finally, using the spatial and temporal components previously estimated as an offset, we estimate the parameters of the spatio-temporal random effect $\zeta(\textbf{u},t)$ and obtain an estimate for the posterior distribution of the intensity function $\lambda(\textbf{u},t)$. In Figure 10, we present the posterior mean of the intensity function for three different days.
Figure 10: Posterior mean of the intensity function for the dates 2020-03-21, 2020-03-22, and 2020-03-23. Black dots are the locations of infected people on these days.
@@ 5.2 Velocities
With the approximation of the intensity function, we can estimate the velocity and its direction using Equations (13), and (15) in the spatial domain. In Figures (11) and (12), we can see the estimate of the absolute value of the temporal derivative in (a), the estimate of the norm of the spatial gradient in (b), and the estimated minimum velocity in (c) for times 2020-03-22 and 2020-03-23 respectively. We presented these two approximations and the velocity (quotient of those two) because we want to show that large values of minimal velocities imply, at the same time, large values of the absolute value of the temporal derivative and lower values of the norm of the spatial gradient.
Specifically, in Figure (11), we can see that in (a) and (b) in the upper right part of the map, there is a large magnitude of the absolute value of the temporal derivative, as well as a lower value of the norm of the gradient. Considering that, it is expected to see a hot spot in the same location as it showed in Figure 11 (c); also, notice the arrows are pointing to the hot spot. This hot spot shows the occurrence of new cases on day 2020-03-22 in the upper right part of Cali, which were not there the day before, as shown in Figure 10. Notice that, as we mentioned before, the arrows point to the hot spot, which means the number of cases in this zone is increasing.
(a) $\left|\partial\lambda(\mathbf{u},t)/\partial t\right|$
(b) $\lVert\nabla\lambda(\textbf{u},t)\rVert$
(c) $s_{\text{min}}(\textbf{u},t)$
Figure 11: Velocity estimation in 2020-03-22.
On the other hand, in Figure 12 (a) and (b), we can see in the lower center of the map a significant value of the absolute value of the temporal derivative, but also a large value of the spatial gradient in the same location. That means that the value of the minimal velocity near this zone will not be as big as we expected; instead, it will be low compared to the other values, as we can see in Figure 12 (c). This happens because despite this zone presents new occurrences in comparison to the day before, as is shown in Figure 10, the density population of Cali is very narrow in that zone, and that will keep the minimal velocity smaller as is shown in Figure 8.
(a) $\left|\partial\lambda(\mathbf{u},t)/\partial t\right|$
(b) $\lVert\nabla\lambda(\textbf{u},t)\rVert$
(c) $s_{\text{min}}(\textbf{u},t)$
Figure 12: Velocity estimation in 2020-03-23.
@@ 6 Conclusions and discussion
In this paper, we present an alternative derivation of the formula for the minimal velocity based on the advection-diffuse equation for a sufficiently smooth function. We use that derivation as a starting point to extend it to the intensity function of a spatio-temporal point process. We use an LGCP model and consider a multiplicative decomposition of the space-time intensity function into spatial, temporal, and residual spatio-temporal components. First, the spatial component representing the expected number of cases can be estimated using a nonparametric estimation by kernels or integrating population data. Second, we proposed a spline model for the temporal component that uses day-of-week, periodic, and polynomial growing effects. Finally, for the spatio-temporal component, we used Matérn covariance structure on space and an independent structure over time. For inference, we used R-INLA and used the estimation of the intensity function to propose an approximation of the minimal velocity using finite differences.
@@
Through a simulation study, we demonstrated that this approach works well, with the approximation by finite differences showing an overall good fit. That happens because the point approximation of the intensity function is smooth enough, which is one of the advantages of approximating an LGCP using basis functions in R-INLA. One of the main contributions of the velocity approximation proposed here is that we are not using a closed form of the covariance structure or a parametric estimation of the spatial or temporal gradients. Here, we proposed a model-free approximation of the velocity that is flexible enough to capture instantenous change in the intensity function between two time steps in the time resolution given by the data, for instance, the change of an infectious disease from one day to the other. We use first-order approximations of the partial derivatives in space and time mainly because of their simplicity. However, for future work, we can improve the approximation by using the smoothing properties of the intensity function to propose second-order approximations of the partial derivatives.
We also presented an application of COVID-19 data in Cali, Colombia, to demonstrate how this methodology works with real data. In this application, we can observe how the minimal velocity could help us understand how the disease spreads from one day to another. This application led us to think of a possible extension of this work. For example, we could model the interaction of more than one infectious disease in the same population, for instance, the evolution of dengue and chikungunya in a specific region [Pavani et al., 2023]. This approach could be extended to understand the marginal spread of each disease, considering their interactions.
@@
In summary, we presented a method to approximate the minimal velocity and direction of disease spread using finite differences of the intensity function modeled with a spatio-temporal process. This is a data-driven approach that uses a flexible model that could be applied in infectious diseases or other spatio-temporal point patterns such as crime data or species locations, taking into account that the covariance structure for these problems is not the same as in contagious diseases.

### v2

@@ 5 COVID-19 spread and velocities in Colombia
In this section, we applied our method to compute minimal velocities to understand the spread of an infectious disease. In particular, we use the household locations of COVID-19-infected individuals in Cali, one of the largest and most populated cities in Colombia, during the COVID-19 pandemic in 2020. Data were provided by the Cali Municipal Public Health Secretary [Municipality_Cali] and have confirmed locations of 208551 infected individuals from 2020-03-01 to 2022-04-25. To begin, the geographical coordinates of the data were converted to UTM coordinates, and the values in the map are presented in meters. To model the LGCP, we used the spatio-temporal SPDE described in (5) with the parameters given by model D. This process involves the construction of a triangulated mesh covering the study region, in this case, the city boundary of Cali, Colombia. Once we have fitted the model, we make predictions in a grid with cells equally spaced 1km apart, which results in 296 cells inside the city limits. For the temporal component, for simplicity, we use the first 80 days. In Figure 4, we present the location of infected people on 4 different days in Cali, Colombia.
Figure 4: Spatial distribution of reported COVID-19 cases in Cali, Colombia, on four selected dates: March 11, March 31, April 20, and May 10, 2020. Each point represents the household location of an infected individual.
@@ 5.1 Intensity function
Following the ideas of Section 3.1, we apply a LGCP model with intensity function decomposed as the product of the spatial $\eta(\textbf{u})$, temporal $\mu(t)$, and a residual spatio-temporal variation $\exp\left(\xi(\textbf{u},t)\right)$ as follows
$\Lambda(\textbf{u},t)=\eta(\textbf{u})\cdot\mu(t)\cdot\exp\left(\xi(\textbf{u},t)\right).$ (20)
In practice, this decomposition requires first computing the spatial and temporal components separately and then using this information to compute the spatio-temporal component. The computation of each component is detailed below.
The spatial component represents the expected number of infected individuals in the city of Cali. To estimate this, we consider as a study region the city boundary of Cali. We use the R-INLA package to construct a triangulated mesh in this region, and compute the expected number of events using the dual mesh, which comprises a series of polygons surrounding each vertex of the original mesh. On the other hand, we obtain the population density in Cali from WorldPop, an open-access tool for population data. In this case, we used a resolution of $100\times 100$ $m$ cells defined in a square region that includes Cali (Figure 5). Then, the expected number of events in each polygon of the dual mesh is computed as the area of the polygon multiplied by the average population of the cells in that polygon [krainski_et_al_2019].
Figure 5: Estimated population density in Cali, Colombia, measured as the number of people in each cell ($100m\times 100m$).
To estimate the temporal component representing the daily expected number of infected people, we use a spline model as explained in Section 3.1.1. The selected basis is related to the structure of the data. It takes into account first the day-of-week effects, periodical effects described by sine and cosine functions called the Fourier basis, and finally an overall rising trend over time captured by a polynomial structure. The selected number of bases in each case was related to the simplicity and explainability of the model. In summary, the model has the following form
@@
Figure 6: Daily COVID-19 cases (blue dots) with the fitted model (solid red line) and its associated prediction interval (shaded pink area) over time. The width of the prediction interval varies, ranging from a lower bound of $1.24$ to an upper bound of $264.3$.
Finally, using the spatial and temporal components previously estimated as an offset, we estimate the parameters of the spatio-temporal random effect $\xi(\textbf{u},t)$ and obtain an estimate for the posterior distribution of the intensity function $\lambda(\textbf{u},t)$ using INLAspacetime and inlabru [lindgren2023inlaspacetime, inlabru]. In Figure 7, we present the posterior mean of the intensity function for three different days.
Figure 7: Posterior mean of the intensity function for March 21, 22, and 23, 2020. Black dots indicate the household locations of infected individuals on each respective day.
@@ 5.2 Velocities
Using the estimated intensity function, we can compute the velocity and its direction within the spatial domain using (15) and (12). Figure 8 illustrates these estimations for March 22, 2020 (a), and March 23, 2020 (b). Specifically, in Figure 8 (a), the upper right region of the map exhibits two zones with high velocity magnitude. The arrows in the upper zone of these two are pointing towards the center, indicating the formation of a hotspot. This pattern suggests the emergence of new cases on March 22, 2020, in the upper right part of Cali, which were not present the previous day, as shown in Figure 7. For the second region, the lack of a clear arrow pattern prevents drawing definitive conclusions.
Additionally, Figure 8 (a) also highlights a high value of the velocity magnitude presented in the left lower region. The arrows in this region, pointing to the left and right of the map, indicate that the number of infected individuals in that region is having a dispersal effect. That means that there was a significant number of cases on the day before that starts to disperse in that region, as it is shown in Figure 7.
(a) Estimated velocity on March 22, 2020.
(b) Estimated velocity on March 23, 2020.
Figure 8: Estimated minimal velocity $s_{\text{min}}(\mathbf{u},t)$ and its direction for March 22 (left) and March 23 (right), 2020.
In Figure 8, the upper right region of the map exhibits two zones with a high velocity magnitude. The arrows in the left zone are pointing towards the center, indicating the formation of a hotspot. This pattern suggests the emergence of new cases on March 23, 2020, in the upper left part of Cali, which were not present the previous day, as shown in Figure 7. On the other hand, the zone on the right lacks a clear pattern, preventing any strong conclusions about its behavior.
We remind readers that this is a data-driven method, and data inaccuracies may affect the model’s accuracy. Such errors may introduce noise in either the magnitude or the direction of the velocity, though typically not in both simultaneously. Therefore, we strongly recommend drawing inferences only when both conditions, high magnitude and clear direction, are present, as the presence of only one is not sufficient.
@@ 6 Conclusions and discussion
In this paper, we present a spatio-temporal modeling approach to estimate the velocities of infectious disease spread. We model the locations and times of people infected using a spatio-temporal log-Gaussian Cox point process, and obtain the velocities using finite differences that approximate the partial derivatives of the intensity function.
@@
Specifically, we present an alternative derivation of the formula for the minimal velocity based on the advection-diffuse equation for a sufficiently smooth function. We use that derivation as a starting point to extend it to the intensity function of a spatio-temporal point process. We use an LGCP model and consider a multiplicative decomposition of the space-time intensity function into spatial, temporal, and residual spatio-temporal components. First, the spatial component representing the expected number of cases can be estimated using a nonparametric estimation by kernels or by integrating population data. Second, we proposed a spline model for the temporal component that uses day-of-week, periodic, and polynomial growing effects. Finally, for the spatio-temporal component, we used a Matérn covariance structure on space and an independent structure over time. For inference, we used INLAspacetime and used the estimation of the intensity function to propose an approximation of the minimal velocity using finite differences.
@@
Through a simulation study, we demonstrated that this approach is effective, with the approximation by finite differences yielding an overall good fit. That happens because the point approximation of the intensity function is smooth enough, which is one of the advantages of approximating an LGCP using basis functions in INLAspacetime. One of the main contributions of the velocity approximation proposed here is that we do not use a parametric estimation of the spatial or temporal gradients. Here, we propose a model-free approximation of the velocity that is flexible enough to capture instantaneous changes in the intensity function between two time steps, given the data’s time resolution, for instance, the change of an infectious disease from one day to the next. We use first-order approximations of the partial derivatives in space and time mainly because of their simplicity. However, for future work, we can improve the approximation by using the smoothing properties of the intensity function to propose second-order approximations of the partial derivatives.
We demonstrated the effectiveness of our method by computing minimal velocities from COVID-19 data in Cali, Colombia, showing how these estimates help to understand disease spread. In this application, we showed that velocity provides richer information about the dynamics of disease propagation than first-order estimators. This insight is especially useful when existing approaches fail to capture directional patterns. In addition, this work could be extended to model the velocities of multiple infectious diseases interacting in the same population, such as dengue and chikungunya, which are vector-borne diseases with common risk factors [pavanietal23]. Our method can also be applied in other fields to study the direction and speed of movement in point patterns – such as species, crimes, or fires – to better understand how their geographic distribution changes over time and how it relates to climate or other factors [moraga20, moraga23].
@@
In summary, we presented a method for approximating the minimal velocity and direction of disease spread using finite differences of the intensity function, modeled with a spatio-temporal process. This is a data-driven approach that uses a flexible model that could be applied in infectious diseases or other spatio-temporal point patterns such as crime data or species locations, taking into account that the covariance structure for these problems is not the same as in contagious diseases.
Appendix A Modeling results
This appendix presents the results of modeling the simulated spatio-temporal point pattern with the intensity function defined in (16). Figures 9, 10, and 11 show, respectively, the intensity function, the absolute value of its temporal derivative, and the norm of its spatial gradient at time points $0.225$, $0.575$, and $0.875$, for $\lambda_{0}=5$. Each figure is organized by columns, with the first column showing the ground truth, and the second and third columns presenting the results from Models C and D, respectively.
Figure 9 compares the true intensity function with the posterior mean from each model. As expected, both models capture the temporal evolution of the intensity. However, Model D (non-separable model) in the third column generally provides a better approximation with smoother contour lines than Model C (separable model) in the second column. This smoothness is particularly important as it reduces noise in derivative computations, improving velocity approximations.
Figure 10 presents the absolute values of the true and approximated temporal derivatives. While the approximations exhibit slightly more noise compared to the intensity estimates, both models successfully capture the overall trend. Notably, the non-separable model (Model D) provides a better fit than the separable model (Model C), particularly at times $0.575$ and $0.875$, reinforcing its advantage in derivative estimation.
Finally, Figure 11 examines the spatial gradient norm for both the true and estimated values. This estimation exhibits even more noise than the temporal derivative approximation due to the additional step of combining spatial derivatives to compute the gradient norm. Despite this, both models capture the overall shape and mode of the true values. Once again, the non-separable model (Model D) demonstrates a better fit and a smoother surface compared to the separable model (Model C), further supporting its effectiveness in velocity estimation.
Intensity function
True Values
Model C
Model D
Time = 0.225
Time = 0.575
Time = 0.875
Figure 9: Intensity estimation for the simulated spatio-temporal point pattern. Columns correspond to the true intensity function (left), the posterior mean obtained using the separable model (Model C, center), and the posterior mean from the non-separable model (Model D, right). Rows represent three key time points in the process: 0.225 (top), 0.575 (middle), and 0.875 (bottom).
Absolute value of temporal derivative of the intensity function
True Values
Model C
Model D
Time = 0.225
Time = 0.575
Time = 0.875
Figure 10: Absolute value of the temporal derivative of the intensity function. Columns correspond to the true derivative (left), the approximation obtained using the separable model (Model C, center), and the approximation from the non-separable model (Model D, right). Rows represent three key time points in the process: 0.225 (top), 0.575 (middle), and 0.875 (bottom).
Norm of the spatial gradient of the intensity function
True Values
Model C
Model D
Time = 0.225
Time = 0.575
Time = 0.875
Figure 11: Norm of the spatial gradient of the intensity function. Columns correspond to the true gradient norm (left), the approximation obtained using the separable model (Model C, center), and the approximation from the non-separable model (Model D, right). Rows represent three key time points in the process: 0.225 (top), 0.575 (middle), and 0.875 (bottom).